OpenStudy (anonymous):

Solve x-6y=19 for y.

hartnn (hartnn):

tried ? subtract x from both sides first

OpenStudy (anonymous):

Well I got x-19/6=y but it doesn't seem right

hartnn (hartnn):

-6y = 19-x 6y = x-19 y= (x-19)/6 did u miss the brackets ?

OpenStudy (anonymous):

Yes! That's what I missed! thanks!

hartnn (hartnn):

welcome ^_^